calculate the momentum of the Titanic (m=4.2*10^7kg) moving at 14 knots (1kt=1.852km/h)
1
Expert's answer
2012-01-13T11:52:22-0500
Let's transform the velocity of Titanic from knots to meters per second:

1kt = 1.852 km/h = 1.852*1000/3600 = 1.852/3.6 m/s;

V = 14 kt = 14 * 1.852/3.6 m/s = 25.928/3.6 m/s;

So, the momentum of the Titanic is

P = mV = 4.2*10^7kg * 25.928/3.6 m/s = 30249.(3) * 10^4 kg*m/s.